# Fast Fourier Transform periodic interpolation method for superposition sums in a periodic unit cell

### v1

@@ Abstract
We propose a Fast Fourier Transform based Periodic Interpolation Method (FFT-PIM), a flexible and computationally efficient approach for computing the scalar potential given by a superposition sum in a unit cell of an infinitely periodic array. Under the same umbrella, FFT-PIM allows computing the potential for 1D, 2D, and 3D periodicities for dynamic and static problems, including problems with and without a periodic phase shift. The computational complexity of the FFT-PIM is of $O(N\log N)$ for $N$ spatially coinciding sources and observer points. The FFT-PIM uses rapidly converging series representations of the Green’s function serving as a kernel in the superposition sum. Based on these representations, the FFT-PIM splits the potential into its near-zone component, which includes a small number of images surrounding the unit cell of interest, and far-zone component, which includes the rest of an infinite number of images. The far-zone component is evaluated by projecting the non-uniform sources onto a sparse uniform grid, performing superposition sums on this sparse grid, and interpolating the potential from the uniform grid to the non-uniform observation points. The near-zone component is evaluated using an FFT-based method, which is adapted to efficiently handle non-uniform source-observer distributions within the periodic unit cell. The FFT-PIM can be used for a broad range of applications, such as periodic problems involving integral equations in computational electromagnetic and acoustic, micromagnetic solvers, and density functional theory solvers.
@@ 1 Introduction
A common category of problem in computational physics is calculating the potential, referred to as periodic scalar potential (PSP), that is generates by a non-uniform distribution of sources arranged in a unit cell that is extended to an infinite periodic array (Fig. 1). Examples of such problems are phased-array antennas, crystals, periodic gratings, periodically modulated waveguides, to name a few [1, 2, 3, 4]. These problems may present challenges for obtaining accurate and rapid solutions. For a standalone unit cell without periodicity, this task can be accomplished via a standard superposition sum involving free-space Green’s function as its kernel. Such sums can be evaluated rapidly in $O(N)$ or $O(N\log N)$ computational cost for $N$ sources and observers using several fast methods, such as fast multipole like method [5, 6], H-matrix method [7], interpolation-based methods [8, 9, 10], and FFT-based methods [11, 12, 13, 14]. However, for the periodic case, the superposition sums involve infinite number of images, which poses additional challenges to the slow converge or even divergence as well as a potentially high computational cost.
@@
Several existing methods can address such periodic problems with certain assumptions and limitations. Periodic Green’s function (PGF) can be defined in terms of an infinite sum over periodic images. In dynamic problems, e.g., in computational electromagnetics, PGF is assumed to be defined with a phase shift or non-vanishing wavenumber [15, 16, 17]. However, for static problems in which no phase shift and wave propagation are present, even in the 1D case, the PGF diverges. In molecular dynamics problems [18], a special form of static PGF assumes uniformly arranged sources [19], such as a cubic lattice. Although various methods have been proposed to address static and dynamic periodic unit cell problems with non-uniform sources, they often focus on specific problem types, such as addressing only 1D or 2D periodicity, addressing only static problems [20], requiring lattice structure or relying on finite difference methods [21]. There is a need in a general approach capable of effectively handling all types of these problems regardless of the periodicity type, distribution of sources, with or without phase shift, and for static and dynamic cases.
In this paper, we proposed a general approach, referred to as Fast Fourier Transform Periodic Interpolation Method (FFT-PIM) that is applicable to all these types of problems under a unified framework, including 1D, 2D, and 3D periodicities for static and dynamic problems with and without periodic phase shifts. The FFT-PIM separates the PSP into a near-zone component involving a small number of images near the periodic unit cell of interest, and a spatially slowly varying far-zone component involving the rest of an infinite number of images. The evaluation of the near-zone PSP component is based on a modified version of the box adaptive integral method [11]. The evaluation of the far-zone PSP component is based on the sparse periodic interpolation method that is adapted to allow handling dynamic and static problems. The computational cost of the FFT-PIM is of $O(N\log N)$ and the memory consumption is of $O(N)$.
@@ 2 Formulation
$u(\mathbf{r}_{m})=\sum_{n=1}^{N}G^{p}(r_{m}-r_{n})q(\mathbf{r}_{n}),$ (1)
@@
$G^{p}_{1D}(\mathbf{r})=$ $\sum_{i_{x}=-\infty}^{\infty}e^{-jk_{x0}(i_{x}L_{x})}G_{0}(\mathbf{r}-i_{x}L_{x}\hat{x}),$ (2a)
$G^{p}_{2D}(\mathbf{r})=$ $\sum_{i_{x}=-\infty}^{\infty}\sum_{i_{y}=-\infty}^{\infty}e^{-j[k_{x0}(i_{x}L_{x})+k_{y0}(i_{y}L_{y})]}G_{0}(\mathbf{r}-i_{x}L_{x}\hat{x}-i_{y}L_{y}\hat{y}),$ (2b)
$G^{p}_{3D}(\mathbf{r})=$ $\sum_{i_{x}=-\infty}^{\infty}\sum_{i_{y}=-\infty}^{\infty}\sum_{i_{z}=-\infty}^{\infty}e^{-j[k_{x0}(i_{x}L_{x})+k_{y0}(i_{y}L_{y})+k_{z0}(i_{z}L_{z})]}G_{0}(\mathbf{r}-i_{x}L_{x}\hat{x}-i_{y}L_{y}\hat{y}-i_{z}L_{z}\hat{z}),$ (2c)
@@
The series of Eq. (2) for the PGF is slowly convergent, which is because of a slow spatial decay of $G_{0}$. An alternative spectral representation of the PGF of Eq. (2) can be based on the Floquet mode expansion [17]:
@@
×$\times\left(\frac{e^{-jk_{zmn}|z|}}{1-e^{-j(k_{zmn}-k_{z0})L_{z}}}-\frac{jk_{zmn}|z|}{1-e^{-j(k_{zmn}+k_{z0})L_{z}}}\right),$ (3c)
where $H_{0}^{(2)}$ is the zeroth order Hankel function of the second kind, $k_{xm}=k_{x0}+2\pi m/L_{x}$, $k_{ym}=k_{y0}+2\pi n/L_{z}$, $k_{\rho m}=(k_{0}^{2}-k_{xm}^{2})^{1/2}$ and $k_{zmn}=(k_{0}^{2}-k_{xm}^{2}-k_{yn}^{2})^{1/2}$, and the square roots are chosen such that their imaginary parts are non-positive. The spectral series of Eq. (2) converges exponentially fast when $(y^{2}+z^{2})^{1/2}$ is non-vanishing and $k_{0},k_{x0},k_{y0},k_{z0}$ are not all vanishing.
A special attention needs to be paid to the case of $k_{0}=k_{x0}=k_{y0}=k_{z0}=0$, which we refer to as static case. The static case is important as it corresponds to several practical problem, such as using periodic extensions to represent infinite 1D, 2D, and 3D domains in various static problems, e.g., to characterize micromagnetic [23] or in molecular dynamics problems. In this static periodic case, the PGF calculated via series of Eq. (2) or Eq. (2) diverges and the PSP for a general source distribution diverges as well. However, a finite PSP can be obtained for the special case of a neutral source, i.e., under the condition
$\sum_{n=1}^{N}q_{n}=0.$ (4)
This neutrality condition corresponds to various physical problems, e.g., when computing the magnetic scalar potential generated by a magnetized objects or electrostatic potential in polarized objects [24]. In this case, the PSP converges and we can neglect any non-relevant components (e.g., constant terms) in the PGF, so that it is convergent as well. The result is that the PGF for the neutral static case can be calculated via:
@@
It can be shown that Eq. (5a) is obtained from Eq. (3a) under the assumption of Eq. (4). In this case, by taking the limit of ($k_{0}\to 0,k_{x0}\to 0$) for all $n\neq 0$, the Hankel function becomes the modified Bessel function and the symmetric ±$\pm n$ terms can be combined into the positive $n$ terms. For the $n=0$ term, the Hankel function exhibits a logarithmic behavior but it is summed up to 00 due to the source neutrality of Eq. (4). The PGF for the 2D and 3D cases can be then obtained by a spatial sum of 1D PGFs. When the sources are arranged on a lattice, the PGF for the periodic static case sometimes is also referred to as the lattice Green’s function [19].
The PGFs converge exponentially fast provided the $(y^{2}+z^{2})^{1/2}$ is not too small. For a desired error of $\epsilon$, the number of terms required Eq. (3a) and Eq. (5a) scales as $L_{x}\log(\epsilon^{-1})/z$ for the 1D case and $L_{x}L_{y}\log^{2}(\epsilon^{-1})/z^{2}$ for the 2D and 3D cases. To demonstrate the convergence, Fig. 2 shows the PGFs and their convergence when using Eq. (2) and Eq. (2) for $L_{x},L_{y},L_{z}=1$, $k_{0}=-1-j$ and $k_{x0}=1-j,k_{y0}=1+j,k_{z0}=-1+j$ in Fig. 2 and the static case in Fig. 2.
Figure 2: Convergence of PGFs, $L_{x},L_{y},L_{z}=1$, $x=y=z=0.5$. (a) The relative error of the sum of the first $m$ terms from Eq. (2), with $k_{0}=-1-j$ and $k_{x0}=1-j,k_{y0}=1+j,k_{z0}=-1+j$, (b) The relative error of the sum of the first $m$ terms from Eq. (2). Exponential convergence is schieved in both cases.
@@ 3.1 Representation of PGF and PSP
Here, the near-zone component $G_{near}^{p}$ is given by Eq. (2) but with the sums over $i_{x}$, $i_{y}$, and $i_{z}$ in the range of $[-i_{xd},i_{xd}]$, $[-i_{yd},i_{yd}]$, and $[-i_{zd},i_{zd}]$, respectively, where $i_{xd},i_{yd},i_{zd}$ are small integer numbers. This components represent the contribution from a small number of images around the zeroth unit cell of interest. Assuming that the periodicities are close to each other, we can use the same value of $i_{d}$ for $i_{xd},i_{yd},i_{yd}$. For the sake of clarity we assume this case in the following discussions. $G_{near}^{p}$ includes the self-term and a few terms from the surrounding images (see Fig. 1 showing the near-zone images for the case of $i_{xd}=i_{yd}=1$ as green boxes). Therefore, $G_{near}^{p}$ may be spatially singular and rapidly varying. The evaluation of $G_{near}^{p}$ involves only a small number of the sum terms.
@@ 3.2 Evaluation of the far-zone PSP
The far-zone PGF and PSP components vary slowly in space, which enables us to calculate the far-zone PGF and PSP on a uniform grid within the observation domain and subsequently interpolate to the required observer points. We construct a sparse uniform grid of observation points $\mathbf{r}_{n(l,p,q)}^{o}={x_{l}^{o},y_{p}^{o},z_{q}^{o}}$} (red grid in Fig. 3) and a spatially shifted uniform grid of source points $\mathbf{r}_{n(l,p,q)}^{s}={x_{l}^{s},y_{p}^{s},z_{q}^{s}}$} (black grid in Fig. 3):
@@
Here, $\omega^{s}(\mathbf{r}_{n^{\prime}}^{s},\mathbf{r}_{n})$ represents the interpolation coefficients from the source grid points $\mathbf{r}_{n^{\prime}}^{s}$ to original non-uniform source points $\mathbf{r}_{n}$ and $\omega^{o}(\mathbf{r}_{m},\mathbf{r}_{m^{\prime}}^{o})$ represents the interpolation coefficients from the observer grid points $\mathbf{r}_{m^{\prime}}^{o}$ to the original non-uniform observer points $\mathbf{r}_{m}$. The number of the source interpolation points $N_{gs}(\mathbf{r}_{n})$ and observer interpolation points $N_{go}(\mathbf{r}_{n})$ represents the number of the grid points that need to be used for the interpolation. For example, we can choose the $q$-th order Lagrange interpolation such that all grid points participate in the interpolation for any of the non-uniform points. For this choice, $N_{q}=(q+1)^{3}$ , $N_{go}=N_{gs}=N_{g}$, and the coefficients $\omega^{o}$ and $\omega^{s}$ are the same for the same points. For a more general case of a grid with $N_{q}$ greater than $N_{go}$ and $N_{gs}$ , the interpolation for each non-uniform point involve only a subset of the grid points, and the coefficients $\omega^{o}$ and $\omega^{s}$ can be represented as sparse matrices. The corresponding sparse matrices are generally transpose versions of each other.
@@
Based on this grid construction in Eq. (3.2) and PSP representation in Eq. (10), we can first pre-compute the coefficients $\omega^{o}$ and $\omega^{s}$, e.g., as a sparse matrix, and pre-compute $G^{p}_{far}(\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s})$ table at the grid points. The computation of $\omega^{o}$ and $\omega^{s}$ involves $O(N)$ operations since the interpolations are spatially local. The computation of $G^{p}_{far}(\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s})$ involves $O(N_{g})$ operations since the grid is uniform and $G^{p}_{far}$ is shift invariant in term of the differences $\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s}$. The computation of $G^{p}_{far}(\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s})$ can be done via the rapidly convergent sums of Eq. (2) and Eq. (2) due to the choice of the source and observer grid in Eq. (3.2) with a spatial separation. The far-zone PSP $u_{far}$ can, then, be evaluated via the following three steps, which are illustrated in Fig. 3.

### v2

@@ Abstract
We propose a Fast Fourier Transform based Periodic Interpolation Method (FFT-PIM), a flexible and computationally efficient approach for computing the scalar potential given by a superposition sum in a unit cell of an infinitely periodic array. Under the same umbrella, FFT-PIM allows computing the potential for 1D, 2D, and 3D periodicities for dynamic problems involving the Helmholtz potential and static problems involving Coulomb potential, including problems with and without a periodic phase shift. The computational complexity of the FFT-PIM is of $O(N\log N)$ for $N$ spatially coinciding sources and observer points. The FFT-PIM uses rapidly converging series representations of the Green’s function serving as a kernel in the superposition sum. Based on these representations, the FFT-PIM splits the potential into its near-zone component, which includes a small number of images surrounding the unit cell of interest, and far-zone component, which includes the rest of an infinite number of images. The far-zone component is evaluated by projecting the non-uniform sources onto a sparse uniform grid, performing superposition sums on this sparse grid, and interpolating the potential from the uniform grid to the non-uniform observation points. The near-zone component is evaluated using an FFT-based method, which is adapted to efficiently handle non-uniform source-observer distributions within the periodic unit cell. The FFT-PIM can be used for a broad range of applications, such as periodic problems involving integral equations for wave propagation in electromagnetics and acoustics, micromagnetic solvers, and density functional theory solvers.
@@ 1 Introduction
A common category of problem in computational physics is calculating the potential, referred to as periodic scalar potential (PSP), that is generates by a non-uniform distribution of sources arranged in a unit cell that is extended to an infinite periodic array (Fig. 1). Examples of such problems are phased-array antennas, crystals, periodic gratings, periodically modulated waveguides, to name a few [1, 2, 3, 4]. These problems may present challenges for obtaining accurate and rapid solutions. For a non-periodic unit cell, this task can be accomplished via a standard superposition sum involving free-space Green’s function as its kernel. Such sums can be evaluated rapidly in $O(N)$ or $O(N\log N)$ computational cost for $N$ sources and observers using several fast methods, such as fast multipole like method [5, 6], H-matrix method [7], interpolation-based methods [8, 9, 10], and FFT-based methods [11, 12, 13, 14]. However, for the periodic case, the superposition sums involve infinite number of images, which poses additional challenges to the slow converge or even divergence as well as a potentially high computational cost.
@@
Several existing methods can address such periodic problems with certain assumptions and limitations[15, 16, 17, 18]. Periodic Green’s function (PGF) can be defined in terms of an infinite sum over periodic images. In dynamic problems, e.g., in computational electromagnetics, PGF is assumed to be defined with a phase shift or non-vanishing wavenumber [19, 20, 21]. However, for static problems with no phase shift, even in the 1D case, the PGF diverges. In molecular dynamics problems [22], a special form of static PGF assumes uniformly arranged sources [23], such as a cubic lattice. Although various methods have been proposed to address static and dynamic periodic unit cell problems with non-uniform sources, they often focus on specific problem types, such as addressing only 1D or 2D periodicity, addressing only static problems [24], requiring lattice structure or relying on finite difference methods [25]. There is a need in a general approach capable of effectively handling all types of these problems regardless of the periodicity type, distribution of sources, with or without phase shift, and for static and dynamic cases.
In this paper, we proposed a general approach, referred to as Fast Fourier Transform Periodic Interpolation Method (FFT-PIM) that is applicable to all these types of problems under a unified framework, including 1D, 2D, and 3D periodicities for static and dynamic problems with and without periodic phase shifts. The FFT-PIM separates the PSP into a near-zone component involving a small number of images near the periodic unit cell of interest, and a spatially slowly varying far-zone component involving the rest of an infinite number of images. The evaluation of the near-zone PSP component is based on the box adaptive integral method [11], which is modified to allow for the rapid computations with multiple near-zone images. The evaluation of the far-zone PSP component is based on the sparse periodic interpolation method [26], which is adapted to allow handling dynamic and static problems with arbitrary phase shifts. The computational cost of the FFT-PIM is of $O(N\log N)$ and the memory consumption is of $O(N)$.
@@ 2 Formulation
$u(\mathbf{r}_{m})=\sum_{n=1}^{N}G^{p}(\mathbf{r}_{m}-\mathbf{r}_{n})q(\mathbf{r}_{n}),$ (1)
@@
$G^{p}_{1D}(\mathbf{r})=$ $\sum_{i_{x}=-\infty}^{\infty}e^{-jk_{x0}(i_{x}L_{x})}G_{0}(\mathbf{r}-i_{x}L_{x}\mathbf{\hat{x}}),$ (2a)
$G^{p}_{2D}(\mathbf{r})=$ $\sum_{i_{x}=-\infty}^{\infty}\sum_{i_{y}=-\infty}^{\infty}e^{-j[k_{x0}(i_{x}L_{x})+k_{y0}(i_{y}L_{y})]}G_{0}(\mathbf{r}-i_{x}L_{x}\mathbf{\hat{x}}-i_{y}L_{y}\mathbf{\hat{y}}),$ (2b)
$G^{p}_{3D}(\mathbf{r})=$ $\sum_{i_{x}=-\infty}^{\infty}\sum_{i_{y}=-\infty}^{\infty}\sum_{i_{z}=-\infty}^{\infty}e^{-j[k_{x0}(i_{x}L_{x})+k_{y0}(i_{y}L_{y})+k_{z0}(i_{z}L_{z})]}G_{0}(\mathbf{r}-i_{x}L_{x}\mathbf{\hat{x}}-i_{y}L_{y}\mathbf{\hat{y}}-i_{z}L_{z}\mathbf{\hat{z}}),$ (2c)
@@
The series of Eq. (2) for the PGF is slowly convergent or even divergent depending on the wavenumbers, which is because of a slow spatial decay of $G_{0}$. An alternative spectral representation of the PGF of Eq. (2) can be based on the Floquet mode expansion [21]:
@@
×$\times\left(e^{-jk_{zmn}|z|}+\frac{e^{-j(k_{zmn}-k_{z0})L_{z}}e^{-jk_{zmn}z}}{1-e^{-j(k_{zmn}-k_{z0})L_{z}}}+\frac{e^{-j(k_{zmn}+k_{z0})L_{z}}e^{jk_{zmn}z}}{1-e^{-j(k_{zmn}+k_{z0})L_{z}}}\right),$ (3c)
where $H_{0}^{(2)}$ is the zeroth order Hankel function of the second kind, $k_{xm}=k_{x0}+2\pi m/L_{x}$, $k_{ym}=k_{y0}+2\pi n/L_{z}$, $k_{\rho m}=(k_{0}^{2}-k_{xm}^{2})^{1/2}$ and $k_{zmn}=(k_{0}^{2}-k_{xm}^{2}-k_{yn}^{2})^{1/2}$, and the square roots are chosen such that their imaginary parts are non-positive. The spectral series of Eq. (3) converges exponentially fast when $(y^{2}+z^{2})^{1/2}$ is non-vanishing and $k_{0},k_{x0},k_{y0},k_{z0}$ are not all vanishing. The series may diverge or not be defined for certain parameter combinations with $k_{zmn}=0$, $k_{\rho m}=0$, which are related to so called Rayleigh-Wood anomalies [28, 29, 30, 31]. For scattering electromagnetic or acoustic problems, such anomalies correspond to the transition between evanescent and propagating Floquet modes. Additional types of Wood anomalies may be of resonant type [31], e.g., when Green’s function is defines in the presence of a layered medium [32]. When such anomalies occur, they may require a special treatment, e.g., Ref. [28] or by defining the phase shift in the complex plane with the integration path deformation [30, 33].
A special attention needs to be paid to the no phase static periodic (NPSP) case where $k_{0}=k_{x0}=k_{y0}=k_{z0}=0$. This case is important as it corresponds to several practical problem, such as using periodic extensions to represent infinite 1D, 2D, and 3D domains in various static problems, e.g., to characterize micromagnetic [34] or in molecular dynamics problems. In this NPSP case, the PGF calculated via series of Eq. (2) or Eq. (3) diverges and the PSP for a general source distribution diverges as well. However, a finite PSP can be obtained for the special case of a neutral source, i.e., under the condition
$\sum_{l=1}^{N}q_{l}=0.$ (4)
This neutrality condition corresponds to various physical problems, e.g., when computing the magnetic scalar potential generated by a magnetized objects or electrostatic potential in polarized objects [35]. In this case, the PSP converges and we can neglect any non-relevant components (e.g., constant terms) in the PGF, so that it is convergent as well. The result is that the PGF for the neutral NPSP case can be calculated via:
@@
It can be shown that Eq. (5a) is obtained from Eq. (3a) under the assumption of Eq. (4). In this case, by taking the limit of ($k_{0}\to 0,k_{x0}\to 0$) for all $n\neq 0$, the Hankel function becomes the modified Bessel function and the symmetric ±$\pm n$ terms can be combined into the positive $n$ terms. For the $n=0$ term, the Hankel function exhibits a logarithmic behavior but it is summed up to 00 due to the source neutrality of Eq. (4). The PGF for the 2D and 3D cases can be then obtained by a spatial sum of 1D PGFs. When the sources are arranged on a lattice, the PGF for the NPSP case sometimes is also referred to as the lattice Green’s function [23].
The PGFs converge exponentially fast provided the $(y^{2}+z^{2})^{1/2}$ is not too small. For a desired error of $\epsilon$, the number of terms required Eq. (3a) and Eq. (5a) scales as $L_{x}\log(\epsilon^{-1})/z$ for the 1D case and $L_{x}L_{y}\log^{2}(\epsilon^{-1})/z^{2}$ for the 2D and 3D cases. To demonstrate the convergence, Fig. 2 shows the PGFs and their convergence when using Eq. (3) and Eq. (5) for $L_{x},L_{y},L_{z}=1$, $k_{0}=-1-j$ and $k_{x0}=1-j,k_{y0}=1+j,k_{z0}=-1+j$ in Fig. 2 and the NPSP case in Fig. 2.
Figure 2: Convergence of PGFs, $L_{x},L_{y},L_{z}=1$, $x=y=z=0.5$. (a) The relative error of the sum of the first $m$ terms from Eq. (3), with $k_{0}=-1-j$ and $k_{x0}=1-j,k_{y0}=1+j,k_{z0}=-1+j$, (b) The relative error of the sum of the first $m$ terms from Eq. (5). Exponential convergence is achieved in both cases.
@@ 3.1 Representation of PGF and PSP
Here, the near-zone component $G_{near}^{p}$ is given by Eq. (2) but with the sums over $i_{x}$, $i_{y}$, and $i_{z}$ in the range of $[-i_{xd},i_{xd}]$, $[-i_{yd},i_{yd}]$, and $[-i_{zd},i_{zd}]$, respectively, where $i_{xd},i_{yd},i_{zd}$ are small integer numbers. This components represent the contribution from a small number of images around the zeroth unit cell of interest. Assuming that the periodic lengths are close to each other, we can use the same value of $i_{d}$ for $i_{xd},i_{yd},i_{yd}$. For the sake of clarity we assume this case in the following discussions. $G_{near}^{p}$ includes the self-term and a few terms from the surrounding images (see Fig. 1 showing the near-zone images for the case of $i_{xd}=i_{yd}=1$ as green boxes). Therefore, $G_{near}^{p}$ may be spatially singular and rapidly varying. The evaluation of $G_{near}^{p}$ involves only a small number of the sum terms.
@@ 3.2 Evaluation of the far-zone PSP
The far-zone PGF and PSP components vary slowly in space, which enables us to calculate the far-zone PGF and PSP on a uniform grid within the observation domain and subsequently interpolate to the required observer points. For a target object of size $D_{x},D_{y},D_{z}$, we construct a sparse uniform grid of observation points $\mathbf{r}_{n(l,p,q)}^{o}={x_{l}^{o},y_{p}^{o},z_{q}^{o}}$} (red grid in Fig. 3) and a spatially shifted uniform grid of source points $\mathbf{r}_{n(l,p,q)}^{s}={x_{l}^{s},y_{p}^{s},z_{q}^{s}}$} (black grid in Fig. 3):
@@
Here, $\omega^{s}(\mathbf{r}_{n^{\prime}}^{s},\mathbf{r}_{n})$ represents the interpolation coefficients from the source grid points $\mathbf{r}_{n^{\prime}}^{s}$ to original non-uniform source points $\mathbf{r}_{n}$ and $\omega^{o}(\mathbf{r}_{m},\mathbf{r}_{m^{\prime}}^{o})$ represents the interpolation coefficients from the observer grid points $\mathbf{r}_{m^{\prime}}^{o}$ to the original non-uniform observer points $\mathbf{r}_{m}$. The number of the source interpolation points $N_{gs}(\mathbf{r}_{n})$ and observer interpolation points $N_{go}(\mathbf{r}_{n})$ represents the number of the grid points that need to be used for the interpolation. For example, we can choose the $q$-th order Lagrange interpolation such that all grid points participate in the interpolation for any of the non-uniform points. For this choice, $N_{q}=(q+1)^{3}$ , $N_{go}=N_{gs}=N_{q}$, and the coefficients $\omega^{o}$ and $\omega^{s}$ are the same for the same points. For a more general case of a grid with $N_{q}$ greater than $N_{go}$ and $N_{gs}$ , the interpolation for each non-uniform point involve only a subset of the grid points, and the coefficients $\omega^{o}$ and $\omega^{s}$ can be represented as sparse matrices. The corresponding sparse matrices are generally transpose versions of each other.
@@
Based on this grid construction in Eq. (8) and PSP representation in Eq. (10), we can first pre-compute the coefficients $\omega^{o}$ and $\omega^{s}$, e.g., as a sparse matrix, and pre-compute $G^{p}_{far}(\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s})$ table at the grid points. The computation of $\omega^{o}$ and $\omega^{s}$ involves $O(N)$ operations since the interpolations are spatially local. The computation of $G^{p}_{far}(\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s})$ involves $O(N_{g})$ operations since the grid is uniform and $G^{p}_{far}$ is shift invariant in term of the differences $\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s}$. The computation of $G^{p}_{far}(\mathbf{r}_{m^{\prime}}^{o}-\mathbf{r}_{n^{\prime}}^{s})$ can be done via the rapidly convergent sums of Eq. (3) and Eq. (5) due to the choice of the source and observer grid in Eq. (8) with a spatial separation. The far-zone PSP $u_{far}$ can, then, be evaluated via the following three steps, which are illustrated in Fig. 3.